# A Game-Theoretic Framework for Distributed Load Balancing: Static and Dynamic Game Models

### v2

@@ 1 Introduction
Distributed and uncoordinated load balancing is challenging when a set of strategic agents compete for a set of shared resources. There is a large body of literature on load balancing, in which the proposed schemes can be either static or dynamic. In static load balancing schemes [1, 2, 3, 4], agents use prior knowledge about the system, whereas dynamic load balancing methods [5, 6, 7, 8, 9] make decisions based on the system’s real-time state.
Load balancing in strategic environments has been studied using game theory in the existing literature, where, in most applications, the system is often modeled as congestion games [10, 11]. In congestion games, each player’s utility is determined by the resources they select and the number of other players choosing the same resources. If all players choose the most beneficial resource, it becomes congested, and all players suffer from the congestion cost. As a result, agents may prefer to choose less profitable resources to avoid congestion.
One of the limitations of the aforementioned models is the assumption that rounds are independent, i.e., the outcome of one round does not affect future rounds [6]. However, in queuing systems, jobs left in queues create a carryover effect, making each round’s average wait time dependent not only on newly assigned jobs but also on residual jobs. Motivated by these limitations and applications in job scheduling and queuing networks, we propose and analyze a game-theoretic framework for load balancing in both static and dynamic settings. Here, jobs are managed by self-interested players who do not coordinate, yet the goal is to achieve decentralized and balanced server loads.
@@ 1.1 Related Work
Load balancing in distributed systems has been extensively studied in the literature. Many of these studies focus on systems with a centralized scheduler [12, 13, 14, 15, 16]. Among the various load balancing approaches, join-the-shortest-queue (JSQ) stands out as one of the most thoroughly researched methods [17, 18, 19, 20, 21, 22, 23]. This approach assigns jobs to the server with the shortest queue length. Join-idle-queue (JIQ) [24, 25, 26] is another load balancing policy where jobs are sent to idle servers, if any, or to a randomly selected server. These methods might not perform well because the queue length is not always a good indicator of the servers’ wait times. Power-of-$d$-choices [15, 27, 28, 29, 30, 31, 32] queries $d$ randomly selected servers and sends the jobs to the least loaded ones. This method has been proven to yield short queues in homogeneous systems but may be unstable in heterogeneous systems.
@@
Studying distributed load balancing through a game-theoretic lens is a natural approach. Despite its potential, relatively few studies have applied game theory to the load balancing problem [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. In [2], the authors model the load balancing problem in heterogeneous systems as a non-cooperative, repeated finite game and identify a structure for the best-response policy in pure strategies. They later introduce a distributed greedy algorithm to implement the best-response policy. However, since their setting is stateless, the proposed algorithm does not necessarily perform well in dynamic games. Additionally, [3] models the load balancing problem as a cooperative game among players. Unlike distributed load balancing algorithms without queue memory [2, 3], which rely on stateless updates, our model explicitly captures the state dynamics and shows that stability and convergence can be guaranteed in polynomial time, even under strategic behavior of the players.
@@
The work [4] studies the distributed load balancing problem using multi-agent reinforcement learning methods. In this approach, each server simultaneously works on multiple jobs, and the efficiency with which the server handles a job depends on its capacity and the number of jobs it processes. The problem is modeled as a stochastic system where each player has access to local information. To evaluate server efficiency, each player maintains an efficiency estimator, a vector summarizing the server’s performance, which also serves as the player’s state. The players’ policies are then formulated as functions of these states. They compare the performance of their algorithm to other existing algorithms, such as non-adaptive algorithms, i.e., algorithms where players use servers in a predetermined manner, while revealing that naive communication between players does not necessarily improve the system’s overall efficiency.
@@
Krishnasamy et al. [9] study a variant of the multi-armed bandit problem with the carryover effect, which is suitable for queuing applications. They consider a bandit setting in which jobs queue for service, and the service rates of the different queues are initially unknown. The authors examine a queuing system with a single queue and multiple servers, where both job arrivals to the queue and the service rates of the servers follow a Bernoulli distribution. At each time slot, only one server can serve the queue, and the objective is to determine the optimal server scheduling policy for each time slot. They further evaluate the performance of different scheduling policies against the no-regret policy [^1] The policy that schedules the optimal server at each time..
@@
More recently, [6] studied the multi-agent version of the queuing system in [9]. Their setting differs from ours in the sense that, in our setting, each server has a separate queue and processes jobs with a fixed service rate. However, [6] describes a system where each player has a queue, and servers process jobs with a fixed time-independent probability. All unprocessed jobs are then sent back to their respective queues. Queues also receive bandit feedback on whether their job was cleared by their chosen server. One concern in such settings is whether the system remains stable, despite the selfish behavior of the learning players. This concern does not arise in settings with a centralized scheduler, as these systems remain stable if the sum of arrival rates is less than the sum of service rates. However, this condition does not guarantee stability in settings with strategic players. In [6], the authors show that if the capacity of the servers is high enough to allow a centralized coordinator to complete all jobs even with double the job arrival rate, and players use no-regret learning algorithms, [^2] A no-regret learning algorithm minimizes the expected difference in queue sizes relative to a strategy that knows the optimal server. then the system remains stable. They later show in [7] that stability can be guaranteed even with an extra capacity of $\frac{e}{e-1}$, where $e\approx 2.718$ is the natural number, but only if players choose strategies that maximize their long-run success rate. Subsequently, [5] extends the model in [6] by allowing servers to hold a single packet for later service if it cannot be served immediately, and by removing the need to prioritize older packets. Compared to no-regret learning in queuing systems [5, 6, 7], which emphasizes learning behavior under adversarial and stochastic arrival processes, our framework adopts a strategic, game-theoretic approach, where agents optimize best responses against evolving queue states. We establish polynomial-time convergence under a deterministic update rule, rather than relying on regret minimization.
@@
Our dynamic game fundamentally differs from repeated congestion games with discounting [44, 45], which assume that each round is independent of the others. First, in repeated homogeneous congestion games, all players sharing a resource receive the same reward, which decreases with increased usage [46]. In contrast, our model allows heterogeneity, where players using the same resource may receive different rewards. Second, we incorporate a carryover effect: queue states evolve across rounds based on their current state and the number of players choosing them. Rather than optimizing discounted payoffs, we focus on long-run system behavior and establish convergence guarantees based on average cost over time.
@@
Finally, our work is also related to [47], where the authors introduce a new class of Markov congestion games, which model non-stationarity in congestion games. They consider the carryover effect, and at each time step $t$ and state $s$, the players participate in a static congestion game. However, the authors show that even with a centralized algorithm, computing a NE remains computationally intractable. Our setting differs in two key aspects: we assume asynchronous policy updates for players and deterministic state transitions. Moreover, in our model, players using the same resource can receive different rewards. Under these assumptions, we show that our algorithm converges to a load-balanced equilibrium state in polynomial time.
@@ 2.2 Dynamic Load Balancing Game
We note that Assumption 1 is the standard stability condition from queuing theory [48] and is crucial in the dynamic setting. Otherwise, one could always consider a worst-case instance where, regardless of any scheduling policy, the length of the waiting jobs in the server queues will grow unbounded.
@@ 3 Theoretical Results for the Static Game
In this section, we analyze the static load balancing game and show that it belongs to the class of exact potential games, for which a pure NE (defined below) is guaranteed to exist [49].
@@
In fact, in exact potential games, best response dynamics—where players iteratively and sequentially redistribute their jobs across servers by best responding to others’ strategies—are known to converge to a pure NE [49]. Thus, we next proceed to characterize the best response dynamics in a closed-form as a solution to a convex program. This closed-form solution will be used in the subsequent section to extend our results to the dynamic game setting. Our main results for the static load balancing game are presented in Lemma 5 and Theorems 2, 7, and 8.
@@
This in view of [49, Lemma 4.4] suggests that the static load balancing game is an exact potential game.
@@
Now suppose player $i$ plays its best-response action to the fixed actions of the other players (including the initial loads on the servers). As in [2], we can write a convex program whose solution gives us the best strategy for player $i$:
@@
The PoA quantifies the inefficiency of Nash equilibria in games due to the selfish behavior of the players, compared to a globally optimal solution [50].
@@ A.2 Proof of Theorem 8
In any pure NE of the static game, each player plays a best response [51], as characterized in Lemma 5. Suppose $a$ is a pure NE. From Eq. (11), for each server $j$ to which player $i$ assigns jobs, we have
@@
Let $a^{\prime}$ be any arbitrary pure Nash equilibrium, and let $a^{*}$ be the action profile that minimizes the social cost. Also define $NE:=SC(a^{\prime})$ and $OPT:=SC(a^{*})$. As in [52], for any $i\in[n]$, we can write

### v3

@@ 1 Introduction
Distributed and uncoordinated load balancing is challenging when a set of strategic agents compete for a set of shared resources. There is a large body of literature on load balancing, in which the proposed schemes can be either static or dynamic. In static load balancing schemes [berenbrink2007distributed, grosu2005noncooperative, grosu2002load, schaerf1994adaptive], agents use prior knowledge about the system, whereas dynamic load balancing methods [abel2025learning, gaitonde2020stability, gaitonde2021virtues, kameda1997comparison, krishnasamy2016regret] make decisions based on the system’s real-time state.
Load balancing in strategic environments has been studied using game theory in the existing literature, where, in most applications, the system is often modeled as congestion games [milchtaich1996congestion, holzman1997strong]. In congestion games, each player’s utility is determined by the resources they select and the number of other players choosing the same resources. If all players choose the most beneficial resource, it becomes congested, and all players suffer from the congestion cost. As a result, agents may prefer to choose less profitable resources to avoid congestion.
One of the limitations of the aforementioned models is the assumption that rounds are independent, i.e., the outcome of one round does not affect future rounds [gaitonde2020stability]. However, in queuing systems, jobs left in queues create a carryover effect, making each round’s average wait time dependent not only on newly assigned jobs but also on residual jobs. Motivated by these limitations and applications in job scheduling and queuing networks, we propose and analyze a game-theoretic framework for load balancing in both static and dynamic settings. Here, jobs are managed by self-interested players who do not coordinate, yet the goal is to achieve decentralized and balanced server loads.
@@ 1.1 Related Work
Load balancing in distributed systems has been extensively studied in the literature. Many of these studies focus on systems with a centralized scheduler [zhou2018degree, hellemans2018power, horvath2019mean, mitzenmacher2001power, lin1992dynamic]. Among the various load balancing approaches, join-the-shortest-queue (JSQ) stands out as one of the most thoroughly researched methods [gupta2007analysis, eschenfeldt2018join, foley2001join, kogias2019r2p2, kogias2020hovercraft, gardner2019smart, winston1977optimality]. This approach assigns jobs to the server with the shortest queue length. Join-idle-queue (JIQ) [lu2011join, mitzenmacher2016analyzing, jennings2015resource] is another load balancing policy where jobs are sent to idle servers, if any, or to a randomly selected server. These methods might not perform well because the queue length is not always a good indicator of the servers’ wait times. Power-of-$d$-choices [mitzenmacher2001power, xie2015power, gardner2017redundancy, ousterhout2013sparrow, roy2015chaos, nasir2015power, zhu2020racksched] queries $d$ randomly selected servers and sends the jobs to the least loaded ones. This method has been proven to yield short queues in homogeneous systems but may be unstable in heterogeneous systems.
@@
Studying distributed load balancing through a game-theoretic lens is a natural approach. Despite its potential, relatively few studies have applied game theory to the load balancing problem [kameda2012optimal, roughgarden2001stackelberg, nash1950bargaining, economides1991multi, economides1990game, orda1993competitive, altman2001routing, korilis1997capacity, koutsoupias1999worst, mavronicolas2001price, roughgarden2002bad]. In [grosu2005noncooperative], the authors model the load balancing problem in heterogeneous systems as a non-cooperative, repeated finite game and identify a structure for the best-response policy in pure strategies. They later introduce a distributed greedy algorithm to implement the best-response policy. However, since their setting is stateless, the proposed algorithm does not necessarily perform well in dynamic games. Additionally, [grosu2002load] models the load balancing problem as a cooperative game among players. Unlike distributed load balancing algorithms without queue memory [grosu2005noncooperative, grosu2002load], which rely on stateless updates, our model explicitly captures the state dynamics and shows that stability and convergence can be guaranteed in polynomial time, even under strategic behavior of the players.
@@
The work [schaerf1994adaptive] studies the distributed load balancing problem using multi-agent reinforcement learning methods. In this approach, each server simultaneously works on multiple jobs, and the efficiency with which the server handles a job depends on its capacity and the number of jobs it processes. The problem is modeled as a stochastic system where each player has access to local information. To evaluate server efficiency, each player maintains an efficiency estimator, a vector summarizing the server’s performance, which also serves as the player’s state. The players’ policies are then formulated as functions of these states. They compare the performance of their algorithm to other existing algorithms, such as non-adaptive algorithms, i.e., algorithms where players use servers in a predetermined manner, while revealing that naive communication between players does not necessarily improve the system’s overall efficiency.
@@
Krishnasamy et al. [krishnasamy2016regret] study a variant of the multi-armed bandit problem with the carryover effect, which is suitable for queuing applications. They consider a bandit setting in which jobs queue for service, and the service rates of the different queues are initially unknown. The authors examine a queuing system with a single queue and multiple servers, where both job arrivals to the queue and the service rates of the servers follow a Bernoulli distribution. At each time slot, only one server can serve the queue, and the objective is to determine the optimal server scheduling policy for each time slot. They further evaluate the performance of different scheduling policies against the no-regret policy [^1] The policy that schedules the optimal server at each time..
@@
More recently, [gaitonde2020stability] studied the multi-agent version of the queuing system in [krishnasamy2016regret]. Their setting differs from ours in the sense that, in our setting, each server has a separate queue and processes jobs with a fixed service rate. However, [gaitonde2020stability] describes a system where each player has a queue, and servers process jobs with a fixed time-independent probability. All unprocessed jobs are then sent back to their respective queues. Queues also receive bandit feedback on whether their job was cleared by their chosen server. One concern in such settings is whether the system remains stable, despite the selfish behavior of the learning players. This concern does not arise in settings with a centralized scheduler, as these systems remain stable if the sum of arrival rates is less than the sum of service rates. However, this condition does not guarantee stability in settings with strategic players. In [gaitonde2020stability], the authors show that if the capacity of the servers is high enough to allow a centralized coordinator to complete all jobs even with double the job arrival rate, and players use no-regret learning algorithms, [^2] A no-regret learning algorithm minimizes the expected difference in queue sizes relative to a strategy that knows the optimal server. then the system remains stable. They later show in [gaitonde2021virtues] that stability can be guaranteed even with an extra capacity of $\frac{e}{e-1}$, where $e\approx 2.718$ is the natural number, but only if players choose strategies that maximize their long-run success rate. Subsequently, [abel2025learning] extends the model in [gaitonde2020stability] by allowing servers to hold a single packet for later service if it cannot be served immediately, and by removing the need to prioritize older packets. Compared to no-regret learning in queuing systems [abel2025learning, gaitonde2020stability, gaitonde2021virtues], which emphasizes learning behavior under adversarial and stochastic arrival processes, our framework adopts a strategic, game-theoretic approach, where agents optimize best responses against evolving queue states. We establish polynomial-time convergence under a deterministic update rule, rather than relying on regret minimization.
@@
Our dynamic game fundamentally differs from repeated congestion games with discounting [krichene2015online, tennenholtz2009learning], which assume that each round is independent of the others. First, in repeated homogeneous congestion games, all players sharing a resource receive the same reward, which decreases with increased usage [rosenthal1973class]. In contrast, our model allows heterogeneity, where players using the same resource may receive different rewards. Second, we incorporate a carryover effect: queue states evolve across rounds based on their current state and the number of players choosing them. Rather than optimizing discounted payoffs, we focus on long-run system behavior and establish convergence guarantees based on average cost over time.
@@
Finally, our work is also related to [cui2022learning], where the authors introduce a new class of Markov congestion games, which model non-stationarity in congestion games. They consider the carryover effect, and at each time step $t$ and state $s$, the players participate in a static congestion game. However, the authors show that even with a centralized algorithm, computing a NE remains computationally intractable. Our setting differs in two key aspects: we assume asynchronous policy updates for players and deterministic state transitions. Moreover, in our model, players using the same resource can receive different rewards. Under these assumptions, we show that our algorithm converges to a load-balanced equilibrium state in polynomial time.
@@ 2.2 Dynamic Load Balancing Game
We note that Assumption 1 is the standard stability condition from queuing theory [kleinrock1975queueing] and is crucial in the dynamic setting. Otherwise, one could always consider a worst-case instance where, regardless of any scheduling policy, the length of the waiting jobs in the server queues will grow unbounded.
@@ 3 Theoretical Results for the Static Game
In this section, we analyze the static load balancing game and show that it belongs to the class of exact potential games, for which a pure NE (defined below) is guaranteed to exist [monderer1996potential].
@@
In fact, in exact potential games, best response dynamics—where players iteratively and sequentially redistribute their jobs across servers by best responding to others’ strategies—are known to converge to a pure NE [monderer1996potential]. Thus, we next proceed to characterize the best response dynamics in a closed-form as a solution to a convex program. This closed-form solution will be used in the subsequent section to extend our results to the dynamic game setting. Our main results for the static load balancing game are presented in Lemma 5 and Theorems 2, 7, and 8.
@@
This in view of [monderer1996potential, Lemma 4.4] suggests that the static load balancing game is an exact potential game.
@@
Now suppose player $i$ plays its best-response action to the fixed actions of the other players (including the initial loads on the servers). As in [grosu2005noncooperative], we can write a convex program whose solution gives us the best strategy for player $i$:
@@
The PoA quantifies the inefficiency of Nash equilibria in games due to the selfish behavior of the players, compared to a globally optimal solution [Koutsoupias1999].
@@ A.2 Proof of Theorem 8
In any pure NE of the static game, each player plays a best response [nash1950equilibrium], as characterized in Lemma 5. Suppose $a$ is a pure NE. From Eq. (11), for each server $j$ to which player $i$ assigns jobs, we have
@@
Let $a^{\prime}$ be any arbitrary pure Nash equilibrium, and let $a^{*}$ be the action profile that minimizes the social cost. Also define $NE:=SC(a^{\prime})$ and $OPT:=SC(a^{*})$. As in [etesami2020smart], for any $i\in[n]$, we can write